### v1

@@ 2.2. $a=4$
If there are at least six blocks of 0’s in $v$ which have lengths at least $\frac{y}{2}$, then we consider the subword of $v$ which contains $n_{1}$ occurrences of 1 and six blocks $0^{\frac{y}{2}}$. Then this word is a subword of the word $v$ and is not a subword of $u$, and its length is $n_{1}+3y+3\leqslant n_{1}+3x$. Then either $n_{1}+3x\leqslant\frac{3}{4}n+4$, in which case we have a proof, or $n_{1}+3x>\frac{3}{4}n+4$. In the latter case we can rewrite the inequality (3) as follows:
$l\leqslant n_{short}-y+1+a=n-n_{1}-4x-y+5=n-(n_{1}+3x)-x-y+5\leqslant\frac{n}{4}-x-y+1,$
which implies the inequality (3’). If there are at most four blocks of 0’s in $v$ which have lengths at least $\frac{y}{2}$, then we can consider the subword of the word $u$ and not of $v$ which has $n_{1}$ occurrences of 1 and five blocks $0^{\frac{y}{2}}$. So, we can get the inequality (3’) similarly to the previous case, where $v$ has at least six blocks of 0’s which have lengths at least $\frac{y}{2}$.
Now we assume that both $u$ and $v$ have five blocks of 0’s with lengths at least $\frac{y}{2}$. We will call them major blocks. Now we prove the following claim:
@@
Let $u$ be a non-special word with $a_{u}=4$. Then there is a word $u^{\prime}$ with $n_{1}$ occurrences of 1 and two blocks $0^{x}$ such that there is at least one and at most two occurrences of $u^{\prime}$ in $u$.
@@
By the conditions of the claim $u_{long}$ has the following form: $u_{long}=0^{x}1^{\alpha}0^{x}1^{\beta}0^{x}1^{\gamma}0^{x}1^{\delta}$ for some positive integers $\alpha,\beta,\gamma,\delta$. Denote $t=min(\alpha,\beta,\gamma,\delta)$; it is obvious that $t<n_{1}/2$. If among the numbers $\alpha,\beta,\gamma,\delta$ at most two are equal to $t$, then there are at most two occurrences of $u^{\prime}=0^{x}1^{t}0^{x}1^{n_{1}-t}$ in $u$. In the other case $u=0^{x}1^{t}0^{x}1^{t}0^{x}1^{t}0^{x}1^{t_{0}}$. Since $u$ is non-special, then $t_{0}\neq t$, $t_{0}\neq 2t$. Since $t_{0}>t$, then $2t<\frac{n_{1}}{2}$. So, there are at most two occurrences of $u^{\prime}=0^{x}1^{2t}0^{x}1^{n_{1}-2t}$ in $u$. The claim is proved.
By claim there is a subword $u^{\prime}$ of $u$ with $n_{1}$ occurrences of 1 and two blocks $0^{x}$ such that there is at least one and at most two subwords of $u$ which are equal to $u^{\prime}$. We let $u_{1}$ and $u_{2}$ denote these subwords, which are equal to $u^{\prime}$ (if $u_{2}$ exists). Then either $v$ does not have a subword which is equal to $u^{\prime}$ or there is a subword $v^{\prime}=u^{\prime}$ of $v$. In the first case $u^{\prime}$ is the subword of the word $u$ and is not a subword of $v$, and $|u^{\prime}|=n_{1}+2x\leqslant n_{1}+\frac{n_{0}}{2}\leqslant\frac{3}{4}n$. In the second case we consider two overlays by 1’s of the word $v$ on $u$ such that $v^{\prime}$ and $u_{1}$ (resp., $v^{\prime}$ and $u_{2}$) coincide. For both overlays we have that since $v\neq u$, then there is a block $0^{s_{1}}$ (resp, $0^{s_{2}}$) in $u$ (probably empty) which is smaller than the corresponding block of 0’s in $v$ (since $s\leqslant y$). We add a block $0^{s_{1}+1}$ ($0^{s_{2}+1}$) to $v^{\prime}$ to the corresponding place. So, we add at most two blocks. Note that for one of the overlays $0^{s_{i}}\neq 0^{y}$; otherwise in both overlays the major blocks of $v$ coincide with the major blocks of $u$. So, there is a unit turn of $u$ which translates the major blocks of $u$ to the major blocks of $u$. Then the subword of $u$ which contains $n_{1}$ occurrences 1’s, all big blocks $0^{x}$ and block $0^{y}$ is equal to $0^{x}1^{t}0^{x}1^{t}0^{x}1^{t}0^{x}1^{t}0^{y}1^{t}$. That is $u_{long}=0^{x}1^{t}0^{x}1^{t}0^{x}1^{t}0^{x}1^{2t}$ is special, a contradiction. So after adding blocks $0^{s_{1}}$ and $0^{s_{2}}$ (one of them contains less than $\frac{y}{2}$ occurrences of 0’s) we get the word $v^{\prime\prime}$ such that $v^{\prime\prime}$ is the subword of the word $v$ and is not a subword of $u$ and $|v^{\prime\prime}|\leqslant n_{1}+2x+y+\frac{y}{2}+2$. Then either lemma is proved, or $n_{1}+2x+y+\frac{y}{2}\geqslant\frac{3}{4}n+2$. But if $n_{1}+2x+y+\frac{y}{2}\geqslant\frac{3}{4}n+2$, then
@@
so $l\leqslant\frac{n}{4}-x-\frac{y}{2}+3$, that is what we needed to get.
3. $a=3$
In this case $u_{fix}$ contains two blocks $0^{x}$. So, we can prove the lemma in this case similarly to the cases 1 and 2, but instead of the case ${\bf 1}$ we will consider the case 1’: $n_{1}+2x+y\leqslant\frac{3}{4}n+3$, and instead of the case ${\bf 2}$ we will consider the case 2’: $n_{1}+2x+y\geqslant\frac{3}{4}n+3$.
In the case 1’ the subword $u_{fix}$ of $u$ is unique and $|u_{fix}|+y+1\leqslant n_{1}+2x+y+1\leqslant\frac{3}{4}n+4$. Then applying Proposition 3.11 to the word $u_{fix}$ we get that there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
In the case 2’, since $a=3$, we can rewrite inequality (3) as follows:
$\begin{split}l\leqslant n_{short}-y+1+a=n-n_{1}-3x-y+a+1=\\ n-(n_{1}+2x+y+1)-x+a\leqslant\\ \frac{n}{4}-x+a-3=\frac{n}{4}-\frac{(a-1)(x-2)}{2}-2.\end{split}$ (3”)
Using (3’’), we can rewrite inequality (4) in the case 2.1.1 as
$|v_{zeros}|+y+1=n_{1}+l+y+1\leqslant n_{1}+\frac{n}{4}-\frac{(a-1)(x-2)}{2}-2+x=$
$=\frac{n}{2}+\frac{n}{4}-\frac{ax-3x-2a+6}{2}=\frac{3}{4}n-\frac{(a-3)(x-2)}{2}\leqslant\frac{3}{4}n+2,$
since $a\geqslant 3$ and $x\geqslant 2$. Using (3’’), we can rewrite inequality (5) in case 2.1.2 as:
$|v_{zeros}|\leqslant\frac{n_{1}+l}{2}+l+2(x-1)+y=\frac{n_{1}+3x+y+(l-4)}{2}+l+\frac{x+y}{2}\leqslant$
@@ 3. $a=3$
since $a\geqslant 3$ and $x\geqslant 2$.
The lemma is proved. ∎
@@
Let $u_{long}$ be periodic and non-special with $a\geqslant 3$. Then there exists a subword of only one of the words $u$ and $v$ of length $\frac{3}{4}n+4$.
Proof. Consider three cases corresponding to the values of $a$.
1. $a\geqslant 5$
In this case $u_{long}=(1^{\alpha_{1}}0^{x}1^{\alpha_{2}}0^{x}\cdots 1^{\alpha_{r}}0^{x})^{t}$, where $rt=a$ and $t\geqslant 2$. So, $r\leqslant a/2$. Since $u_{long}$ is not special, then $r\geqslant 2$. Consider the word $u_{1}=(1^{\alpha_{1}}0^{x}1^{\alpha_{2}}0^{x}\cdots 1^{\alpha_{r}}0^{x})1^{r(t-1)}$. Then either $u_{1}$ is not a subword of $v$, or there is a subword $v_{1}=u_{1}$ of $v$. In the first case $u_{1}$ is the subword of the word $u$ and is not a subword of $v$, and $|u_{1}|=n_{1}+\frac{ax}{t}\leqslant n_{1}+\frac{n_{0}}{2}\leqslant\frac{3}{4}n$. In the second case there are at most $t$ overlays by 1’s of $v$ on $u$ such that $v_{1}$ coincide with equal to $v_{1}$ subword. For each such overlay there is a block $0^{s}$ $(s\leqslant y)$ in $u$ which is smaller than the corresponding block in $v$. We add $0^{s+1}$ to $v_{1}$ to the corresponding place; let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$. Since $y+1\leqslant x$, then
$|v_{1}^{\prime}|\leqslant n_{1}+rx+\frac{a}{r}(y+1)\leqslant n_{1}+rx+(\frac{a}{r}-1)x+(y+1).$
Consider the function $f(r)=r+\frac{a}{r}-1$. It is easy to see that for $a\geqslant 5$, in the interval $22\leqslant r\leqslant a/$ 2 we have $f(r)\leqslant\max(f(2),f(\frac{a}{2}))=\frac{a}{2}+1$. Then
@@ 1. $a\geqslant 5$
So if $n_{1}+\frac{ax}{2}+x+y\leqslant\frac{3}{4}n+3$, then we can take $v^{\prime}$ as a required subword.
Assume that $n_{1}+\frac{ax}{2}+x+y\geqslant\frac{3}{4}n+3$. Similarly to case 2 of Lemma 4.2 we can obtain the inequality (3): $l\leqslant\frac{n}{4}-\frac{(a-2)(x-2)}{2}$. Now we proceed with the proof similarly to the case $2$ from Lemma 4.2.
If $u_{zeros}$ is not periodic, then the proof is similar to case 2.1.1 from Lemma 4.2. Assume that $u_{zeros}$ is periodic. That is, $u_{zeros}=(1^{\alpha_{1}}01^{\alpha_{2}}0\ldots 1^{\alpha_{p}}0)^{\frac{l}{p}}$. Consider the word $u_{zeros}^{\prime}=(1^{\alpha_{1}}01^{\alpha_{2}}0\ldots 1^{\alpha_{p}}0)(10)^{l-p}$. Similarly to the case 2.1.2 of Lemma 4.2, the word $u_{zeros}^{\prime}$ contains $l$ occurrences of 0 and at most $\frac{n_{1}+l}{2}$ occurrences of 1. We add to $u_{zeros}^{\prime}$ all blocks $0^{x}$ from $u_{1}$ (recall that $u_{1}=(1^{\alpha_{1}}0^{x}1^{\alpha_{2}}0^{x}\ldots 1^{\alpha_{r}}0^{x})1^{r(t-1)}$). We let $u_{zeros}^{\prime\prime}$ denote the obtained subword of $u$. So either $u_{zeros}^{\prime\prime}$ is not a subword of $v$ and $|u_{zeros}^{\prime\prime}|\leqslant\frac{3}{4}n+4$ (it will follow from further inequalities), or there is a subword $v^{\prime}=u_{zeros}^{\prime\prime}$ of $v$. In the last case there are at most $t$ overlays by 1’s of $v$ on $u$ such that $v^{\prime}$ coincides with a subword which equal to $v^{\prime}$. For each of these overlays there is a block $0^{s}$ $(s\leqslant y)$ in $u$ which is less than the corresponding block in $v$. We add $0^{s+1}$ to $v_{1}$ to the corresponding place. We let $v^{\prime\prime}$ denote the obtained subword of $v$. The word $v^{\prime\prime}$ is a subword of the word $v$ and is not a subword of $u$ and it has at most $\frac{n_{1}+l}{2}$ occurrences of 1, one zero from each block of 0’s, $rx$ occurrences of 0 from big blocks $0^{x}$ and $t$ added blocks each of which contains at most $y+1$ occurrences of 0’s. Since $rt=a$ and $r,t\geqslant 2$, then $r+t\leqslant\frac{a+2}{2}+1$. Since $r+t\leqslant\frac{a+2}{2}+1$, $y\leqslant x-1$, then
@@
The last inequality can be proved similarly to the inequality (5) from Lemma 4.2.
2. $a=4$
Since $u_{long}$ is not special, then $u_{long}=0^{x}1^{\alpha}0^{x}1^{\beta}0^{x}1^{\alpha}0^{x}1^{\beta}$, $\alpha\neq\beta$. This case is proved in the same way as the case 2.2 from Lemma 4.2.
3. $a=3$
This case is impossible since if $u_{long}$ is periodic with three blocks $0^{x}$, then $u_{long}$ is special.
Lemma is proved. ∎
@@ 3. $a=3$
Suppose that one of the following two conditions holds:
$\bullet$ both words are special and $a=b=2$.
$\bullet$ $u$ is a non-special word with at most two blocks $0^{x}$ and $v$ is either special or contains at most two blocks $0^{x}$.
Then there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.
@@
Recall that $0^{y_{u}}=0^{y}$ and $0^{y_{v}}$ are the second largest blocks of 0’s in $u$ and $v$, respectively. Without loss of generality we may assume that $y\geqslant y_{v}$. We now consider the overlay by 1’s of the word $u$ on $v$ such that the blocks $0^{x}$ coincide; we let $\pi$ denote this overlay. If for this overlay $\pi$ there are two neighboring 1’s such that there are 0’s between them in only one of the words $u$ and $v$, then the word $v_{1}=0^{y+1}1^{\alpha}01^{n_{1}-\alpha}$ (or $u_{1}=0^{y_{v}+1}1^{\alpha}01^{n_{1}-\alpha}$) for some $\alpha$ is a subword of only one of the words $u$ and $v$. Moreover, the length of this word at most $n_{1}+y+2\leqslant n_{1}+(x+y)/2+2\leqslant n_{1}+n_{0}/2+2\leqslant\frac{3}{4}n+2$. So, in this case we have a required subword.
@@
Otherwise, for overlay $\pi$ for each place where $u$ has a block of 0’s there is a block of 0’s in $v$ and vice versa. In this case, there are two blocks $0^{\beta}$ and $0^{\gamma}$ in $u$ such that $0^{\beta}$ is bigger than the corresponding block in $v$ and $0^{\gamma}$ is smaller than the corresponding block in $v$. Without loss of generality we assume that $\beta\leqslant\gamma$ (the case $\beta>\gamma$ is similar). Then the word $u_{1}=0^{y+1}1^{\alpha}0^{\beta}1^{n_{1}-\alpha}$ (for some $\alpha$) is a subword of the word $u$ and is not a subword of $v$ and $|u_{1}|=n_{1}+(y+1)+\beta=n_{1}+n_{0}-x-(n_{short}-y-\beta)+1\leqslant n-x-l+4$, since $n_{short}-y-\beta\geqslant l-3$. So, if $n-x-l+4\leqslant\frac{3}{4}n+4$, then we have a required subword. If $n-x-l+4>\frac{3}{4}n+4$, then we get $x+l<\frac{n}{4}$. Since for the overlay $\pi$ for each place where $u$ has a block of 0’s there is a block of 0’s in $v$ and vice versa, then $u_{2}=0^{y+1}1(01)^{\alpha-1}0^{\beta}1(01)^{n_{1}-\alpha-1}$ is a subword of the word $u$ and is not a subword of $v$ and $|u_{2}|=2l+y+\beta-1<2l+2x<\frac{n}{2}$. We proved the lemma in case 1.
2. $(a,b)=(2,1)$.
@@
In this case the word $u_{1}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ is a subword of the word $u$ and is not a subword of $v$ and $|u_{1}|=n-n_{short}$. If $n-n_{short}\leqslant\frac{3}{4}n$, then we have the required subword, so it remains to consider the case $n_{short}<\frac{n}{4}-4$. This inequality implies that $l\leqslant n_{short}-y+3<\frac{n}{4}-y-1$. The word $u_{2}=0^{x}1(01)^{\alpha_{1}-1}0^{x}1(01)^{\alpha_{2}-1}$ is a subword of the word $u$ and is not a subword of $v$, and $|u_{2}|=2l+2x-2$. If $2l+2x-2\leqslant\frac{3}{4}n$, then we again have a required subword, so it remains to consider the case $42l+2x-2>\frac{3}{4}n+$ 4. The latter inequality implies that $42(\frac{n}{4}-y-1)+2x-2>\frac{3}{4}n+$ 4, which can be rewritten as $82x-2y\geqslant\frac{n}{4}+$ 8. Now consider the subword $v_{3}=1^{n_{1}}0^{y+1}$ of $v$. There are two overlays by 1’s of $v$ on $u$ such that $v_{3}$ coincides with an equal subword. For both overlays we find blocks $0^{s_{1}}$ and $0^{s_{2}}$ which are less than the corresponding blocks in $v$ and we add to $v_{3}$ blocks $0^{s_{1}+1}$ and $0^{s_{2}+1}$ to the corresponding places; we let $v_{3}^{\prime}$ denote the obtained subword of $v$. The word $v_{3}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$, and
@@
since $n_{1}+y+2x\leqslant n$ and $82(y-x)\leqslant-\frac{n}{4}-$ 8. We proved the lemma in case 2.
3. $(a,b)=(2,2)$.

### v2

@@ 4.2 Lemmas constituting the proof
If there are at least six blocks of 0’s in $v$ which have lengths at least $\frac{y}{2}$, then we consider the subword of $v$ which contains $n_{1}$ occurrences of 1 and six blocks $0^{\frac{y}{2}}$. Then this word is a subword of the word $v$ and is not a subword of $u$, and its length is $n_{1}+3y+3\leqslant n_{1}+3x$. Then either $n_{1}+3x\leqslant\frac{3}{4}n+4$, in which case we conclude, or $n_{1}+3x>\frac{3}{4}n+4$. In the latter case we can rewrite inequality (3) as follows:
$l\leqslant n_{\operatorname{short}}-y+1+a=n-n_{1}-4x-y+5=n-(n_{1}+3x)-x-y+5\leqslant\frac{n}{4}-x-y+1,$
which implies inequality ($3^{\prime}$). If there are at most four blocks of 0’s in $v$ which have lengths at least $\frac{y}{2}$, then we can consider the subword of the word $u$ and not of $v$ which has $n_{1}$ occurrences of 1 and five blocks $0^{\frac{y}{2}}$. So, we can get ($3^{\prime}$) similarly to the previous case, where $v$ has at least six blocks of 0’s which have lengths at least $\frac{y}{2}$.
Now we assume that both $u$ and $v$ have five blocks of 0’s with lengths at least $\frac{y}{2}$. We call them major blocks. Now we prove the following claim:
@@
Let $u$ be not a special word with $a_{u}=4$. Then there is a word $u^{\prime}$ with $n_{1}$ occurrences of 1 and two blocks $0^{x}$ such that there is at least one and are at most two occurrences of $u^{\prime}$ in $u$.
@@
[…]
By the claim there is a subword $u^{\prime}$ of $u$ with $n_{1}$ occurrences of 1 and two blocks $0^{x}$ such that there is at least one and there are at most two subwords of $u$ which are equal to $u^{\prime}$. We let $u_{1}$ and $u_{2}$ denote these subwords, which are equal to $u^{\prime}$ (if $u_{2}$ exists). Then either $v$ does not have a subword which is equal to $u^{\prime}$ or there is a subword $v^{\prime}=u^{\prime}$ of $v$. In the first case $u^{\prime}$ is the subword of the word $u$ and is not a subword of $v$, and $|u^{\prime}|=n_{1}+2x\leqslant n_{1}+\frac{n_{0}}{2}\leqslant\frac{3}{4}n$. In the second case we consider two 1-overlays of the word $v$ on $u$ such that $v^{\prime}$ and $u_{1}$ (resp., $v^{\prime}$ and $u_{2}$) coincide. For both overlays we have that since $v\neq u$, there is a block $0^{s_{1}}$ (resp, $0^{s_{2}}$) in $u$ (possibly empty) which is smaller than the corresponding block of 0’s in $v$ (since $s\leqslant y$). We add a block $0^{s_{1}+1}$ ($0^{s_{2}+1}$) to $v^{\prime}$ in the corresponding place. So, we add at most two blocks. Note that for one of the overlays $0^{s_{i}}\neq 0^{y}$; otherwise in both overlays the major blocks of $v$ coincide with the major blocks of $u$. So, there is a turn of $u$ which translates the major blocks of $u$ to the major blocks of $u$. Then the subword of $u$ which contains $n_{1}$ occurrences of 1, all big blocks $0^{x}$ and block $0^{y}$ is equal to $0^{x}1^{t}0^{x}1^{t}0^{x}1^{t}0^{x}1^{t}0^{y}1^{t}$. That is $u_{\operatorname{long}}=0^{x}1^{t}0^{x}1^{t}0^{x}1^{t}0^{x}1^{2t}$ is special; a contradiction. So after adding blocks $0^{s_{1}}$ and $0^{s_{2}}$ (one of them contains less than $\frac{y}{2}$ occurrences of 0) we get the word $v^{\prime\prime}$ such that $v^{\prime\prime}$ is a subword of the word $v$ and is not a subword of $u$ and $|v^{\prime\prime}|\leqslant n_{1}+2x+y+\frac{y}{2}+2$. Then either $v^{\prime\prime}$ is a distinguishing subword of a desired length, or $n_{1}+2x+y+\frac{y}{2}\geqslant\frac{3}{4}n+2$. But if $n_{1}+2x+y+\frac{y}{2}\geqslant\frac{3}{4}n+2$, then
@@
so $l\leqslant\frac{n}{4}-x-\frac{y}{2}+3$, which is what we needed to get.
Case 3. $a=3$.
In this case $u_{\operatorname{fix}}$ contains two blocks $0^{x}$. So, we can prove the lemma in this case similarly to Cases 1 and 2, but instead of Case 1 we consider Case $\bf 1^{\prime}$}: $n_{1}+2x+y\leqslant\frac{3}{4}n+3$, and instead of Case 2 we consider Case $\bf 2^{\prime}$: $n_{1}+2x+y\geqslant\frac{3}{4}n+3$.
In Case $\bf 1^{\prime}$} the subword $u_{\operatorname{fix}}$ of $u$ is unioccurrent and $|u_{\operatorname{fix}}|+y+1\leqslant n_{1}+2x+y+1\leqslant\frac{3}{4}n+4$. Then applying Proposition 3.11 to the word $u_{\operatorname{fix}}$, we get that there exists a distinguishing subword of length at most $\frac{3}{4}n+4$ for $u$ and $v$.
In Case $\bf 2^{\prime}$}, since $a=3$, we can rewrite inequality (3) as follows:
$l\leqslant n_{\operatorname{short}}-y+1+a=n-n_{1}-3x-y+a+1=$
$n-(n_{1}+2x+y+1)-x+a\leqslant\\ \frac{n}{4}-x+a-3=\frac{n}{4}-\frac{(a-1)(x-2)}{2}-2.$ ($3^{\prime\prime}$)
Using ($3^{\prime\prime}$), we can rewrite inequality (4) in Case 2.1.1 as
$|v_{\operatorname{zeros}}|+y+1=n_{1}+l+y+1\leqslant n_{1}+\frac{n}{4}-\frac{(a-1)(x-2)}{2}-2+x=$
$\frac{n}{2}+\frac{n}{4}-\frac{ax-3x-2a+6}{2}=\frac{3}{4}n-\frac{(a-3)(x-2)}{2}\leqslant\frac{3}{4}n+2,$
since $a\geqslant 3$ and $x\geqslant 2$. Using ($3^{\prime\prime}$), we can rewrite inequality (5) in Case 2.1.2 as:
$|v_{\operatorname{zeros}}|\leqslant\frac{n_{1}+l}{2}+l+2(x-1)+y=\frac{n_{1}+3x+y+(l-4)}{2}+l+\frac{x+y}{2}\leqslant$
@@
since $a\geqslant 3$ and $x\geqslant 2$. ∎
@@
Let $u_{\operatorname{long}}$ be periodic and not special with $a\geqslant 3$. Then there exists a distinguishing subword of length at most $\frac{3}{4}n+4$.
Proof. Since for $a=3$ we have that $u_{long}$ is special as it is periodic with exactly three blocks $0^{x}$, we only have to look at the cases when $a\geq 4$.
Case 1. $a\geqslant 5$.
In this case $u_{\operatorname{long}}=(1^{\alpha_{1}}0^{x}1^{\alpha_{2}}0^{x}\cdots 1^{\alpha_{r}}0^{x})^{t}$ for some $r$ and $t$ such that $rt=a$ and $t\geqslant 2$. So, $r\leqslant a/2$. Since $u_{\operatorname{long}}$ is not special, we have $r\geqslant 2$. Consider the word $u_{1}=(1^{\alpha_{1}}0^{x}1^{\alpha_{2}}0^{x}\cdots 1^{\alpha_{r}}0^{x})1^{r(t-1)}$. Then either $u_{1}$ is not a subword of $v$, or there is a subword $v_{1}=u_{1}$ of $v$. In the first case $u_{1}$ is a subword of the word $u$ and is not a subword of $v$, and $|u_{1}|=n_{1}+\frac{ax}{t}\leqslant n_{1}+\frac{n_{0}}{2}\leqslant\frac{3}{4}n$. In the second case there are at most $t$ 1-overlays of $v$ on $u$ such that $v_{1}$ coincides with a an occurrence of a subword equal to $v_{1}$. For each such 1-overlay there is a block $0^{s}$ $(s\leqslant y)$ in $u$ which is smaller than the corresponding block in $v$. We add $0^{s+1}$ to $v_{1}$ to the corresponding place; let $v_{1}^{\prime}$ denote the obtained subword of $v$. The word $v_{1}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$. Since $y+1\leqslant x$, we have
$|v_{1}^{\prime}|\leqslant n_{1}+rx+\frac{a}{r}(y+1)\leqslant n_{1}+rx+\left(\frac{a}{r}-1\right)x+(y+1).$
Consider the function $f(r)=r+\frac{a}{r}-1$. It is easy to see that for $a\geqslant 5$, in the interval $2\leqslant r\leqslant a/2$ we have $f(r)\leqslant\max(f(2),f(\frac{a}{2}))=\frac{a}{2}+1$. Then
@@
So, if $n_{1}+\frac{ax}{2}+x+y\leqslant\frac{3}{4}n+3$, then we can take $v^{\prime}$ as a required subword.
Assume that $n_{1}+\frac{ax}{2}+x+y\geqslant\frac{3}{4}n+3$. Similarly to Case 2 from Lemma 4.2 we can obtain inequality (3): $l\leqslant\frac{n}{4}-\frac{(a-2)(x-2)}{2}$. Now we proceed with the proof similarly to Case 2 from Lemma 4.2.
If $u_{\operatorname{zeros}}$ is not periodic, then the proof is similar to Case 2.1.1 from Lemma 4.2. Assume that $u_{\operatorname{zeros}}$ is periodic. That is, $u_{\operatorname{zeros}}=(1^{\alpha_{1}}01^{\alpha_{2}}0\cdots 1^{\alpha_{p}}0)^{\frac{l}{p}}$, for some $p\geq 2$. Consider the word $u_{\operatorname{zeros}}^{\prime}=(1^{\alpha_{1}}01^{\alpha_{2}}0\ldots 1^{\alpha_{p}}0)(10)^{l-p}$. Similarly to Case 2.1.2 from Lemma 4.2, the word $u_{\operatorname{zeros}}^{\prime}$ contains $l$ occurrences of 0 and at most $\frac{n_{1}+l}{2}$ occurrences of 1. We add to $u_{\operatorname{zeros}}^{\prime}$ all blocks $0^{x}$ from $u_{1}$ (recall that $u_{1}=(1^{\alpha_{1}}0^{x}1^{\alpha_{2}}0^{x}\ldots 1^{\alpha_{r}}0^{x})1^{r(t-1)}$). We let $u_{\operatorname{zeros}}^{\prime\prime}$ denote the obtained subword of $u$. So either $u_{\operatorname{zeros}}^{\prime\prime}$ is not a subword of $v$ and $|u_{\operatorname{zeros}}^{\prime\prime}|\leqslant\frac{3}{4}n+4$ (we will prove this inequality later), or there is a subword $v^{\prime}=u_{\operatorname{zeros}}^{\prime\prime}$ of $v$. In the latter case there are at most $t$ 1-overlays of $v$ on $u$ such that $v^{\prime}$ coincides with an occurrence of a subword equal to $v^{\prime}$. For each of these overlays there is a block $0^{s}$ $(s\leqslant y)$ in $u$ which is smaller than the corresponding block in $v$. We add $0^{s+1}$ to $v^{\prime}$ to the corresponding place. We let $v^{\prime\prime}$ denote the obtained subword of $v$. The word $v^{\prime\prime}$ is a subword of the word $v$ and is not a subword of $u$ and it has at most $\frac{n_{1}+l}{2}$ occurrences of 1, one zero from each block of 0’s, $rx$ occurrences of 0 from big blocks $0^{x}$ and $t$ added blocks each of which contains at most $y+1$ occurrences of 0. Since $rt=a$ and $r,t\geqslant 2$, we have $r+t\leqslant\frac{a+2}{2}+1$. Since $r+t\leqslant\frac{a+2}{2}+1$, $y\leqslant x-1$, we have
@@
The last inequality can be proved similarly to inequality (5) from Lemma 4.2.
Case 2. $a=4$.
Since $u_{\operatorname{long}}$ is not special, we have $u_{\operatorname{long}}=0^{x}1^{\alpha}0^{x}1^{\beta}0^{x}1^{\alpha}0^{x}1^{\beta}$, $\alpha\neq\beta$. This case is proved in the same way as Case 2.2 from Lemma 4.2. ∎
@@
Suppose $u_{\operatorname{long}}$ is not a special word with at most two blocks $0^{x}$ and $v_{\operatorname{long}}$ is either special or contains at most two blocks $0^{x}$ (or vice versa). Then there is a distinguishing subword of length at most $\frac{3}{4}n+4$.
@@
Recall that $0^{y_{u}}=0^{y}$ and $0^{y_{v}}$ are the second largest blocks of 0’s in $u$ and $v$, respectively. Without loss of generality we may assume that $y\geqslant y_{v}$. We now consider a 1-overlay of $u$ on $v$ such that the blocks $0^{x}$ coincide; we let $\pi$ denote this overlay. If for this 1-overlay $\pi$ there are two neighboring 1’s such that there are 0’s between them in only one of the words $u$ and $v$, then the word $v_{1}=0^{y+1}1^{\alpha}01^{n_{1}-\alpha}$ (or $u_{1}=0^{y_{v}+1}1^{\alpha}01^{n_{1}-\alpha}$) for some $\alpha$ is a distinguishing subword for the words $u$ and $v$. Moreover, the length of this word is at most $n_{1}+y+2\leqslant n_{1}+(x+y)/2+2\leqslant n_{1}+n_{0}/2+2\leqslant\frac{3}{4}n+2$. So, in this case we have a required subword.
@@
Otherwise, for a 1-overlay $\pi$ for each place where $u$ has a block of 0’s there is a block of 0’s in $v$ and vice versa. In this case, there are two blocks $0^{\beta}$ and $0^{\gamma}$ in $u$ such that $0^{\beta}$ is bigger than the corresponding block in $v$ and $0^{\gamma}$ is smaller than the corresponding block in $v$. Without loss of generality we assume that $\beta\leqslant\gamma$ (the case $\beta>\gamma$ is similar). Then for some integer $\alpha$ the word $u_{1}=0^{y+1}1^{\alpha}0^{\beta}1^{n_{1}-\alpha}$ is a subword of the word $u$ and is not a subword of $v$ and $|u_{1}|=n_{1}+(y+1)+\beta=n_{1}+n_{0}-x-(n_{\operatorname{short}}-y-\beta)+1\leqslant n-x-l+4$, since $n_{\operatorname{short}}-y-\beta\geqslant l-3$. So, if $n-x-l+4\leqslant\frac{3}{4}n+4$, then we have a required subword. If $n-x-l+4>\frac{3}{4}n+4$, then we get $x+l<\frac{n}{4}$. Since for a 1-overlay $\pi$ for each place where $u$ has a block of 0’s there is a block of 0’s in $v$ and vice versa, we have that $u_{2}=0^{y+1}1(01)^{\alpha-1}0^{\beta}1(01)^{l-\alpha-1}$ is a subword of the word $u$ and is not a subword of $v$ and $|u_{2}|=2l+y+\beta-1<2l+2x<\frac{n}{2}$. We proved the lemma in Case 1.
Case 2. $(a,b)=(2,1)$.
@@
In this case the word $u_{1}=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}$ is a subword of the word $u$ and is not a subword of $v$ and $|u_{1}|=n-n_{\operatorname{short}}$. If $n-n_{\operatorname{short}}\leqslant\frac{3}{4}n$, then we have a required subword, so it remains to consider the case $n_{\operatorname{short}}<\frac{n}{4}-4$. This inequality implies that $l\leqslant n_{\operatorname{short}}-y+3<\frac{n}{4}-y-1$. The word $u_{2}=0^{x}1(01)^{\alpha_{1}-1}0^{x}1(01)^{\alpha_{2}-1}$ is a subword of the word $u$ and is not a subword of $v$, and $|u_{2}|=2l+2x-2$. If $2l+2x-2\leqslant\frac{3}{4}n$, then we again have a required subword, so it remains to consider the case $2l+2x-2>\frac{3}{4}n+4$. The latter inequality implies that $2(\frac{n}{4}-y-1)+2x-2>\frac{3}{4}n+4$, which can be rewritten as $2x-2y\geqslant\frac{n}{4}+8$. Now consider the subword $v_{3}=1^{n_{1}}0^{y+1}$ of $v$. There are two 1-overlays of $v$ on $u$ such that $v_{3}$ coincides with an occurrence of an equal subword. For both overlays we find blocks $0^{s_{1}}$ and $0^{s_{2}}$ which are shorter than the corresponding blocks in $v$, and we add to $v_{3}$ blocks $0^{s_{1}+1}$ and $0^{s_{2}+1}$ to the corresponding places; we let $v_{3}^{\prime}$ denote the obtained subword of $v$. The word $v_{3}^{\prime}$ is a subword of the word $v$ and is not a subword of $u$, and
@@
since $n_{1}+y+2x\leqslant n$ and $2(y-x)\leqslant-\frac{n}{4}-8$. This completes the proof in Case 2.
Case 3. $(a,b)=(2,2)$.